OpenStudy (anonymous):

domain in interval notation of g(x) = 3x/x^2-4

OpenStudy (anonymous):

I assume it's \[g(x) = \frac{ 3x }{ x^2 - 4 }\]First, find the values of x for which this function is undefined. That can be done by setting the denominator to 0.\[x^2 - 4 = 0\]\[(x - 2)(x + 2) = 0\]So, x cannot be 2 and -2.Thus, the domain is \[x \in (\infty , -2) \cup (2 , \infty)\]

OpenStudy (anonymous):

thank you!